# Improving the Cosmological Constraints by Inferring the Formation Channel of Extreme-mass-ratio Inspirals

### v1

@@ 7 Conclusion
In this work, we have investigated the potential of using EMRIs as dark sirens to constrain the cosmological parameters as well as the formation channel of EMRIs. We adopted 11 population models presented in Babak et al. (2017) to simulate the detected EMRI GW event catalogs.
@@
(i) To constrain the cosmological parameters, we derive the redshift probability distributions of detected EMRIs by matching their spatial localization error volumes with galaxy survey catalogs, and then evaluate the posterior probability distributions of the cosmological parameters via a Bayesian analytical framework which combines the luminosity distance and redshift information. Our results show that precise constraints on the Hubble-Lemaître constant $H_{0}$ and the dark energy EoS parameter $w_{0}$ can be obtained from EMRIs. Under the fiducial population model M1, the precision of $H_{0}$ can reach about $8.1\%$ for TianQin, about $4.4\%$ for TianQin I+II, and about $1.9\%$ for the TianQin+LISA network; and the precision of $w_{0}$ can reach about $37\%$ for TianQin, about $12\%$ for TianQIn I+II, and about $6.5\%$ for the TianQin+LISA network. Additionally, under the most optimistic EMRI population model, TianQin can constrain $H_{0}$ and $w_{0}$ to a precision better than about $3\%$ and about $10\%$, and the TianQin+LISA network can constrain $H_{0}$ and $w_{0}$ to a precision better than about $1\%$ and about $5\%$, respectively.
@@
(ii) To constrain the formation channel of EMRIs, we adopt an AGN luminosity function to randomly generate the numbers of candidate host AGNs of the detected EMRIs, and then use Poisson distributions to statistically test the spatial correlation between the EMRIs and AGNs. Under the fiducial population model, M1, we mainly find that the minimum number $N_{{\rm GW},3\sigma}^{\rm threshold}$ of the EMRIs required to establish the EMRI-AGN correlation decreases with increasing $f_{\rm agn}$. If $f_{\rm agn}=1$, for the three detector configurations of TianQin, TianQin I+II, and TianQin+LISA, the values of $N_{{\rm GW},3\sigma}^{\rm threshold}$ respectively are about $465$, $196$, and $26$ when the cosmological parameters are fixed. Under the condition that the cosmological parameters vary freely within the priors, the values of $N_{{\rm GW},3\sigma}^{\rm threshold}$ become significantly larger, all by a factor of about three or a bit larger. For other population models, $N_{{\rm GW},3\sigma}^{\rm threshold}$ is similar. Moreover, after we confirm that a large fraction of EMRIs do indeed originate in AGNs, the combination of the EMRIs and AGN catalogs can significantly improve the constraints on the cosmological parameters.
@@
The cosmological constraints reported in this work appear to be somewhat weaker than those of LISA alone reported in Laghi et al. (2021) (also see MacLeod & Hogan, 2008). For example, under the M1, M5, and M6 models, LISA with a four year mission is expected to constrain $H_{0}$ and $w_{0}$ to higher precisions than TianQin or TianQin I+II. This is mainly because the most sensitive band of TianQin is slightly higher than that of LISA, which makes the EMRI detection rate of TianQin much lower than LISA (Babak et al., 2017; Fan et al., 2020). However, we would like to point out that currently there is large uncertainty in the mass function of MBHs especially at the low mass end (Greene & Ho, 2007; Barausse, 2012; Gallo & Sesana, 2019; He et al., 2024). The GW signals radiated by the EMRIs that are formed by lighter MBHs, will have higher frequencies and are more easily detected by TianQin, thus enhancing the total EMRI detection rate of TianQin.
@@
Our analytical results of constraining the EMRI-AGN correlation demonstrate for the first time the prospects of using space-borne GW detectors to statistically infer the formation channel of EMRIs. Information about some intrinsic parameters of EMRI, such as the eccentricity at plunge (Pan et al., 2021) and the spin of MBH (Burke et al., 2020), can help us trace the formation channel of the EMRI, but an unbiased extraction of the intrinsic parameters relies heavily on an accurate EMRI GW waveform model (Chua et al., 2017; Amaro-Seoane, 2018; Torres-Orjuela et al., 2021). The effectiveness of the statistical framework proposed in this work mainly relies on the spatial localization accuracies of EMRIs and the completenesses of AGN catalogs. Since the localization accuracies are mainly determined by the SNRs of the EMRIs and the completenesses of the AGN catalogs are mainly determined by the depths of galaxy surveys, the accuracy of the waveform model is not as essential in our framework as in other studies. Therefore, our statistical framework can serve as an important supplementary means of inferring the formation channel of EMRIs.
@@ Acknowledgements
The authors would like to thank Linhua Jiang, Yuming Fu, Zhenwei Lyu and Tieguang Zi for very helpful discussions, and thank the Theoretical Astrophysical Observatory (TAO; https://tao.asvo.org.au/tao/) for providing access to the download of the mock galaxy catalogs. L.-G. Zhu is funded by China Postdoctoral Science Foundation (Grant No. 2023M740113), X. Chen is supported by the National Key Research and Development Program of China (Grant No. 2021YFC2203002) and the National Natural Science Foundation of China (Grant No. 11991053), Y.-M. Hu is supported by the the Natural Science Foundation of China (Grants No. 12173104 and No. 12261131504), and Guangdong Major Project of Basic and Applied Basic Research (Grant No. 2019B030302001), and J.-d. Zhang is supported by the Guangdong Basic and Applied Basic Research Foundation (Grant No. 2023A1515030116). This work also supported by Key Laboratory of TianQin Project (Sun Yat-sen University), Ministry of Education. The computations in this work was performed on the High Performance Computing Platform of the Centre for Life Science, Peking University.

### v2

@@ 7 Conclusion
In this work, we have investigated the potential of using EMRIs as dark sirens to constrain the cosmological parameters as well as infer the formation channel of EMRIs. We adopted 11 population models presented in Babak et al. (2017) to generate catalogs of EMRIs which are detectable by different space GW detectors or detector networks..
@@
(i) To constrain the cosmological parameters, we derived the redshift probability distributions of detected EMRIs by matching their spatial localization error volumes with galaxy survey catalogs, and then evaluated the posterior probability distributions of the cosmological parameters via a Bayesian analytical framework. Our results show that precise constraints on the Hubble-Lemaître constant $H_{0}$ and the dark energy EoS parameter $w_{0}$ can be obtained from EMRIs. Under the fiducial population model M1, the precision of $H_{0}$ can reach about $8.1\%$ for TianQin, about $4.4\%$ for TianQin I+II, and about $1.9\%$ for the TianQin+LISA network; and the precision of $w_{0}$ can reach about $37\%$ for TianQin, about $12\%$ for TianQIn I+II, and about $6.5\%$ for the TianQin+LISA network. Additionally, under the most optimistic EMRI population model, TianQin can constrain $H_{0}$ and $w_{0}$ to a precision better than about $3\%$ and about $10\%$, and the TianQin+LISA network can constrain $H_{0}$ and $w_{0}$ to a precision better than about $1\%$ and about $5\%$, respectively.
@@
(ii) To constrain the formation channel of EMRIs, we adopted an AGN luminosity function to self-consistently generate the numbers of candidate host AGNs for the detected EMRIs, and then used the Poisson distribution to statistically test the spatial correlation between EMRIs and AGNs. We mainly found that the minimum number $N_{{\rm GW},3\sigma}^{\rm threshold}$ of the EMRIs required to establish the EMRI-AGN correlation decreases with increasing $f_{\rm agn}$. If $f_{\rm agn}=1$ and the cosmological parameters are fixed, the values of $N_{{\rm GW},3\sigma}^{\rm threshold}$ for the three detector configurations, namely TianQin, TianQin I+II, and TianQin+LISA, are about $1300$, $450$, and $30$ respectively. Under the condition that the cosmological parameters vary freely within the priors, the values of $N_{{\rm GW},3\sigma}^{\rm threshold}$ become significantly larger, by a factor of about three or larger. However, in our simulated cases where the correlation between EMRIs and AGNs can be statistically confirmed, a combination of the EMRI and AGN catalogs can significantly improve the constraints on the cosmological parameters.
@@
The cosmological constraints reported in this work appear to be somewhat weaker than those of LISA alone reported in Laghi et al. (2021) (also see MacLeod & Hogan, 2008). For example, under the M1, M5, and M6 models, LISA with a four year mission is expected to constrain $H_{0}$ and $w_{0}$ to higher precisions than TianQin or TianQin I+II. This is mainly because the most sensitive band of TianQin is slightly higher than that of LISA, which makes the EMRI detection rate of TianQin much lower than LISA (Babak et al., 2017; Fan et al., 2020). However, we would like to point out that currently there is large uncertainty in the mass function of MBHs especially at the low mass end (Greene & Ho, 2007; Barausse, 2012; Gallo & Sesana, 2019; He et al., 2024). The GW signals radiated by the EMRIs that are formed by lighter MBHs, will have higher frequencies and are more easily detected by TianQin.
@@
Our results also demonstrate the prospect of using space-borne GW detectors to statistically infer the formation channel of EMRIs. It is previously known that the intrinsic parameters of an EMRI, such as the eccentricity at plunge (Pan & Yang, 2021; Pan et al., 2021) and the spin of the MBH (Burke et al., 2020), can help us infer the formation channel of the EMRI, but an unbiased extraction of the intrinsic parameters relies heavily on an accurate EMRI waveform model (Chua et al., 2017; Amaro-Seoane, 2018; Torres-Orjuela et al., 2021). Now the effectiveness of the statistical framework proposed in this work mainly relies on the spatial localization accuracies of EMRIs and the completeness of AGN catalogs. Since the localization accuracies are mainly determined by the SNRs of the EMRIs and the completeness of the AGN catalogs is mainly determined by the depths of AGN surveys, the demand of accurate waveform model subsides to a secondary role. In this sense, our statistical framework could serve as a useful complementary means of inferring the formation channel of EMRIs.
@@ Acknowledgements
The authors would like to thank Linhua Jiang, Yuming Fu, Zhenwei Lyu, Zhen Pan and Tieguang Zi for very helpful discussions, thank the anonymous referee for many helpful comments, and thank the Theoretical Astrophysical Observatory (TAO; https://tao.asvo.org.au/tao/) for providing access to the download of the mock galaxy catalogs. This work is supported by the National Key Research and Development Program of China (Grant No. 2021YFC2203002) and the National Natural Science Foundation of China (Grant No. 11991053). L.-G. Zhu is funded by China Postdoctoral Science Foundation (Grant No. 2023M740113), H.-M. Fan is supported by the Hebei Natural Science Foundation (Grant No. A2023201041), Y.-M. Hu is supported by the Natural Science Foundation of China (Grants No. 12173104 and No. 12261131504), and Guangdong Major Project of Basic and Applied Basic Research (Grant No. 2019B030302001), and J.-d. Zhang is supported by the Guangdong Basic and Applied Basic Research Foundation (Grant No. 2023A1515030116). This work is also supported by Key Laboratory of TianQin Project (Sun Yat-sen University), Ministry of Education. The computations in this work was performed on the High Performance Computing Platform of the Centre for Life Science, Peking University.
Appendix A Distributions of the likelihood ratio in the null and alternative hypotheses
In this appendix we illustrate a deviation of the distribution of the likelihood ratio from the background distribution as $f_{\rm agn}$, the fraction of EMRIs originated in AGNs, increases. We assume that $100$ EMRIs are detected by the TianQin+LISA network, which is expected by most population models for five years of observation, regardless of whether EMRIs originate predominantly in normal galaxies or AGNs (Babak et al., 2017; Fan et al., 2020; Pan et al., 2021). The background distributions $P_{\rm bg}(\lambda)$ of the likelihood ratio are shown in Figure 10 as the dashed lines, while the probability distributions $P(\lambda)$ of the “signal” likelihood ratio (derived from detected EMRIs) are represented by solid lines (we wish to clarify that the signal likelihood ratio $\lambda$ is a deterministic function of $f_{\rm agn}$ in actual GW detections). Both distributions $P_{\rm bg}(\lambda)$ and $P(\lambda)$ are derived from tens of thousands of independently repeated simulations. We can see that the variation of $P(\lambda)$ is more significant than $P_{\rm bg}(\lambda)$ as $f_{\rm agn}$ increases. This discrepancy arises because the variation of $P(\lambda)$ is caused by a combination of the increase in the parameter $f_{\rm agn}$ and the resulting increasing number of candidate host AGNs $\{N_{{\rm agn},i}\}$, whereas the variation of $P_{\rm bg}(\lambda)$ solely stems from the change in the parameter $f_{\rm agn}$.
Figure 10: Typical background distribution $P_{\rm bg}(\lambda)$ (dashed line) and probability distribution $P(\lambda)$ (solid line) of the likelihood ratio $\lambda$ in Equation (18) for 100 EMRIs detected by TianQin+LISA, assuming our fiducial cosmological parameters. The black, blue, cyan, and red colors correspond to four fractions of $f_{\rm agn}=0.1,\!~{}0.3,\!~{}0.6,\!~{}1.0$, respectively.